# WorldLLM: Improving LLMs’ world modeling using curiosity-driven theory-making

### v2

@@ 1 Introduction
Large Language Models (LLMs) possess broad knowledge about the world, but leveraging this knowledge for precise dynamics modeling remains challenging (vafa_evaluating_2024). While LLMs can engage in general reasoning, they struggle to make accurate predictions in specific domains with structured observations and dynamics, such as physics simulations or video games. This limitation stems from the gap between their general capabilities and the need for grounded, domain-specific understanding (carta_grounding_2023; cross_hypothetical_2024; bender_climbing_2020).
@@
The challenge of learning world models has been extensively studied in model-based reinforcement learning (RL), as depicted in moerland_model-based_2022. Traditional approaches focus on learning forward dynamics models through direct interactions, typically using neural networks to predict next observations given current states and actions. Recent work has shown promise in using LLMs for world modeling by fine-tuning them on embodied experiences (xiang_language_2023). However, such an approach is particularly computationally costly given the large amount of interaction data required as well as the cost of gradient-based finetuning of LLMs. Contrary to this, LLMs are known to possess general knowledge that can be adapted using careful prompting (zhang_language-guided_2024).
@@
In a parallel line of work, theory-based RL takes inspiration from humans’ inductive reasoning and intuitive theory-making to suggest focusing on inferring structured and compositional world models (e.g. programs) (tsividis_human-level_2021). Current approaches either generate programs to capture environment dynamics directly (tsividis_human-level_2021; tang_worldcoder_2024; li_automated_2024) or use natural language rules as an intermediate representation (piriyakulkij_doing_2024; wang_hypothesis_2023). Leveraging the world model’s structure as well as human-provided prior distributions over possible world models, theory-based methods are usually more sample-efficient than gradient-based ones. More importantly, the resulting world model often shows stronger generalization abilities. However, contrary to gradient-based approaches that usually do not require any expert-given knowledge, theory-based RL methods mostly rely on hand-crafted theory spaces. Finally, one major challenge in learning a world model lies in the data collection. Indeed, collecting transitions that cover the space of possible outcomes in the environment is key for learning an accurate world model. Notably, both classic model-based RL approach and theory-based RL methods agree on the importance of active data collection focusing on evidence with low likelihood under the current model (schmidhuber_curious_1991; pathak_curiosity-driven_2017; burda_exploration_2018; piriyakulkij_doing_2024).
@@
In this paper, we present WorldLLM, a framework for autonomous improvement of an LLM’s world modeling abilities. Our approach combines 1) probabilistic theory induction to produce hypotheses that are given in our LLM’s prompt to improve its predictions and 2) curiosity-driven RL to explore the environment and collect transitions poorly predicted by the current hypotheses. Formally, our LLM’s world model is the conditional probability $P(s_{t+1}|s_{t},a_{t},H)$, where $s_{t}$ represents a state, $a_{t}$ an action, and $H$ a set of natural language hypothesized theories. This probability is computed by the LLM by giving it $s_{t}$, $a_{t}$, and $H$ in its prompt and taking the probability of $s_{t+1}$ to follow this prompt. Our key insight is that natural language theories can help ground an LLM’s broad knowledge into precise predictive power by providing domain-specific rules (zhang_language-guided_2024). Our approach consists of three interacting components: (1) our LLM that computes $P(s_{t+1}|s_{t},a_{t},H)$ by conditioning its predictions on both a state-action pair and the current hypotheses, (2) a theory generator that updates natural language hypotheses using Bayesian inference, and (3) a curiosity-driven reinforcement learning agent trained to collect evidence against the current hypotheses. Inspired by how humans, from children to scientists (piaget_construction_1954; gopnik_words_1997), actively update their internal world model by performing experiments, our agent’s exploration provides new evidence for hypothesis refinement, creating a virtuous cycle of improvement.
@@ 2.1 Problem statement
In such a setting, we study how to rapidly learn an LLM-based parametric model $\mathbb{P}^{LLM}$ of the transition function $\mathcal{T}$ (also called forward model) given a budget of experiences $T$ allowed to an agent. This parametric model outputs a probability distribution over all possible next states given a state-action pair. The natural approach from the RL literature to improve this forward model uses gradient-based updates (ha_world_2018; hafner_dream_2020; schrittwieser_mastering_2020). However, the scale of modern LLMs makes such an approach particularly costly. LLMs’ natural in-context learning abilities open up an alternative approach: providing natural language hypotheses in the prompt to improve their predictions. This method is less costly, known as more sample efficient (le_scao_how_2021), and provides human-interpretable updates. Consequently, the parameter we seek to optimize is a set of hypotheses $\mathcal{H}\subseteq\mathcal{V}^{N}$ that, when given as input to the forward model (in addition to state-action pairs), maximizes the likelihood of the experienced next states.
@@ 2.2 WorldLLM
To improve the Statistician’s prediction as a forward model, the Scientist’s role is to search through the space of possible sets of hypotheses. To efficiently explore this space, we take inspiration from (wang_hypothesis_2023; piriyakulkij_doing_2024) and use a Bayesian inference framework. In particular, we rely on another LLM $\mathbb{P}^{Sc}$ as the proposal distribution which generates a candidate set of hypotheses following the distribution $\mathbb{P}^{Sc}(\hat{H}_{t}^{i}|\mathcal{D}_{t},\hat{H}_{t})$ given a data set of collected transitions $\mathcal{D}_{t}$ and the current best hypotheses candidate $\hat{H_{t}}$. We assume a uniform prior (i.e., no particular prior) over the space of hypotheses. For this paper, we chose the Metropolis algorithm, a variant of the Metropolis-Hastings algorithm, which iteratively generates candidate hypotheses for $n_{s}teps$. The full algorithm is displayed in Algorithm 1 and more details are provided in Appendix B.
@@
Then, we introduce RL-based Experimenters. We propose to study three curiosity-based intrinsic rewards derived from $\mathbb{P}^{St}$ (oudeyer_what_2007). First, we introduce RL-LogP, where the Experimenter uses the prediction error (schmidhuber_curious_1991; pathak_curiosity-driven_2017; burda_large-scale_2018) as reward by receiving $r=-log\mathbb{P}^{St}(s^{\prime}|s,a,H_{t})$ for each collected transition $(s,a,s^{\prime})$. While usually efficient, this method is also known to suffer in stochastic environments where it fails to separate epistemic from aleatoric uncertainty (burda_exploration_2018). Second, we propose RL-ALP, which leverages information gain over the prediction error (a form of Learning Progress) (oudeyer_what_2007; lopes_exploration_2012; schmidhuber_curious_1991) and rewards transitions on which the forward model improves. In particular, we use the absolute value of Learning Progress, which tracks both progress and forgetting in the forward model (oudeyer_what_2007; baranes_r-iac_2009):
@@
Finally, directly tracking Learning Progress at the transition level often leads to noisy estimations. We also propose RL-ALPEXP, a variant of RL-ALP which partitions the transition space into subspaces and computes ALP over each subspace, stabilizing the estimations. While existing works studied how to automatically infer such partitions (baranes_active_2013), we here rely on hand-designed ones, which correspond to the partitions used by oracles. Further details are provided in Appendix A.
@@ 3.1 Experimental setup
To evaluate WorldLLM’s efficiency, we use the Playground-Text environment from (gaven2024sacglamimprovingonlinerl), a textual environment which produces textual observations and receives textual commands. This environment generates a room that contains multiple objects the agent can interact with by grasping or releasing them. The agent can combine two objects, if an interaction is possible, by releasing one onto the other, as illustrated in Figure 2 b. For instance, releasing water on a carrot seed transforms them into grown carrot. In our experiments, we only consider 4 types of objects: Water, Plants (seed or grown) that require water to grow, Small Herbivores (baby or grown) that require a grown plant to grow, and Big Herbivores (baby or grown) that require 2 grown plants to grow. We evaluate WorldLLM’s ability to discover, predict, and explain how to grow Plants, Small Herbivores, and Big Herbivores. We provide more details on the environment in Appendix A. For analysis, we gather the possible transitions into 6 different types: Standing when moving on an object, Holding 1 when grasping an object, Holding 2 when grasping two objects at the same time, Grow Plant when releasing water on a seed, Grow S. Herbivore when releasing a grown plant on a baby small herbivore and Grow B. Herbivore when releasing two grown plants on a baby big herbivore. These types are the ones used by RL-ALPEXP.
@@ 4.1 Building forward models
Building forward models is a long-established research direction (ha_world_2018; moerland_model-based_2022). In addition to producing more interpretable agents, forward models were particularly used in RL to improve sample efficiency, allowing the agent to train on imagined transitions or perform multi-step planning (hafner_dream_2020; hafner_mastering_2022; hafner_mastering_2024; schrittwieser_mastering_2020). Parallel to this, learning a forward model has been used as an intrinsic reward to foster RL policies’ exploration (schmidhuber_possibility_1991; pathak_curiosity-driven_2017; burda_large-scale_2018; oudeyer_what_2007). WorldLLM employs a similar curiosity-driven scheme where our RL-based Experimenters use reward signals from prior work to promote exploration and collect transitions that help improve the forward model. However, our method differs in the way the model is improved: instead of using a gradient-based approach to fine-tune the LLM-based forward model, as done in (xiang_language_2023), WorldLLM searches for natural language theories that improve the forward model when given in the prompt. While recent work studied how a forward model could leverage external textual information such as a manual explaining the environment (zhang_language-guided_2024; dainese_reader_2023), our approach iteratively generates and refines its own explanations.
@@ 4.2 Inductive reasoning for world modeling
The importance of theories in shaping one’s representations about the world was argued as central, in both children development and adults (piaget_construction_1954; gopnik_scientist_1996) Building on this, the literature of theory-based RL has been advocating for augmenting artificial agents with inductive reasoning (tsividis_human-level_2021). In particular, (tsividis_human-level_2021; tang_worldcoder_2024; li_automated_2024) proposed approaches based on LLMs that generate programs representing the theories about an environment’s dynamics. The use of programs allowed the authors to execute and compare how the proposed theories match empirical evidence. Parallel to this, (piriyakulkij_doing_2024; wang_hypothesis_2023) proposed to generate theories using natural language. While the former constrained the generated theories to programs that can be easily evaluated in their environment, the latter used the generated theories to condition a program generator to simulate the environment’s dynamics. WorldLLM also builds on the theory-based RL literature and closely relates to (piriyakulkij_doing_2024). Notably, WorldLLM also leverages Bayesian inference with an LLM as the proposal distribution to generate the theories. However, the major difference lies in how the produced theories are used: while (piriyakulkij_doing_2024) constrain the format of theories such that they can be directly tested in the environment, our work focuses on giving these theories in another LLM’s prompt to improve its forward modeling capabilities. In other words, our work strongly differs in its perspective of producing theories that improve our LLM’s world modeling abilities instead of producing correct theories. Nonetheless, the Bayesian inference framework we used in our Scientist is largely inspired by (piriyakulkij_doing_2024). Both our work and theirs also consider an active learning paradigm to choose which experiments to perform for theories’ refinement. While their approach computes exact Information Gain of all possible experiments, ours relies on curiosity-driven RL agents and Information Gain approximates as rewards to overcome the intractability of exact Information Gain in most environments.
@@ 5 Discussion
Despite promising results, our initial implementation of WorldLLM presents several limitations. First, the Playground-Text environment is structurally simple, and the natural ordering created by the most challenging transitions being longer and harder to predict is not expected in more complex environments—potentially reducing the effectiveness of RL-LogP. Evaluating WorldLLM in more complex environments is therefore a critical direction for future work (ying_assessing_2025).
@@
In addition to scaling the LLM, we argue that the emergence of such abstract hypotheses may result from other modifications to our implementation. A key advantage of Bayesian inference in this context is the ability to incorporate priors, which constrain the theory space. While a length-based prior might encourage abstract hypotheses, we believe richer priors—such as “error maps” that represent structured patterns of predictive failure, as described by schulz_chapter_2012—may be necessary to better replicate human-like theory refinement.
@@
Finally, our exploration strategy relied solely on simple curiosity-driven signals focused on poorly predicted transitions. Other intrinsic motivation mechanisms have been studied in both humans and artificial agents. A particularly important one is autotelic learning, where learners generate, select, and attempt to master self-defined goals (steels_autotelic_2004; colas_autotelic_2022). Studying the use of such more advanced curiosity signals could significantly improve the quality of collected transitions.
@@ Appendix A The Playground-Text Environment
We use the Playground-Text environment from (gaven2024sacglamimprovingonlinerl), itself being a textual adaptation of the Playground environment introduced by (colas_language_2020). This textual environment returns, at each time step, a description of its state and possible actions. In our experiments, we used a modified version of Playground-Text where the current state still describes the whole environment’s state, but the next state only describes how the performed action changed the environment (see Figure 6).

### v3

@@ 1 Introduction
Large Language Models (LLMs) possess broad knowledge about the world, but leveraging this knowledge for precise dynamics modeling remains challenging (Vafa et al., 2024). While LLMs can engage in general reasoning, they struggle to make accurate predictions in specific domains with structured observations and dynamics, such as physics simulations or video games. This limitation stems from the gap between their general capabilities and the need for grounded, domain-specific understanding (Carta et al., 2023; Cross et al., 2025; Bender and Koller, 2020).
@@
The challenge of learning world models has been extensively studied in model-based reinforcement learning (RL), as depicted in Moerland et al. (2023). Traditional approaches focus on learning forward dynamics models through direct interactions, typically using neural networks to predict next observations given current states and actions. Recent work has shown promise in using LLMs for world modeling by fine-tuning them on embodied experiences (Xiang et al., 2023). However, such an approach is particularly computationally costly given the large amount of interaction data required as well as the cost of gradient-based finetuning of LLMs. Contrary to this, LLMs are known to possess general knowledge that can be adapted using careful prompting (Zhang et al., 2024).
@@
In a parallel line of work, theory-based RL takes inspiration from humans’ inductive reasoning and intuitive theory-making to suggest focusing on inferring structured and compositional world models (e.g. programs) (Tsividis et al., 2021). Current approaches either generate programs to capture environment dynamics directly (Tsividis et al., 2021; Tang et al., 2024; Li et al., 2024) or use natural language rules as an intermediate representation (Piriyakulkij et al., 2024; Wang et al., 2024). Leveraging the world model’s structure as well as human-provided prior distributions over possible world models, theory-based methods are usually more sample-efficient than gradient-based ones. More importantly, the resulting world model often shows stronger generalization abilities. However, contrary to gradient-based approaches that usually do not require any expert-given knowledge, theory-based RL methods mostly rely on hand-crafted theory spaces. Finally, one major challenge in learning a world model lies in the data collection. Indeed, collecting transitions that cover the space of possible outcomes in the environment is key for learning an accurate world model. Notably, both classic model-based RL approach and theory-based RL methods agree on the importance of active data collection focusing on evidence with low likelihood under the current model (Schmidhuber, 1991a; Pathak et al., 2017; Burda et al., 2019b; Piriyakulkij et al., 2024).
@@
In this paper, we present WorldLLM, a framework for autonomous improvement of an LLM’s world modeling abilities. Our approach combines 1) probabilistic theory induction to produce hypotheses that are given in our LLM’s prompt to improve its predictions and 2) curiosity-driven RL to explore the environment and collect transitions poorly predicted by the current hypotheses. Formally, our LLM’s world model is the conditional probability $P(s_{t+1}|s_{t},a_{t},H)$, where $s_{t}$ represents a state, $a_{t}$ an action, and $H$ a set of natural language hypothesized theories. This probability is computed by the LLM by giving it $s_{t}$, $a_{t}$, and $H$ in its prompt and taking the probability of $s_{t+1}$ to follow this prompt. Our key insight is that natural language theories can help ground an LLM’s broad knowledge into precise predictive power by providing domain-specific rules (Zhang et al., 2024). Our approach consists of three interacting components: (1) our LLM that computes $P(s_{t+1}|s_{t},a_{t},H)$ by conditioning its predictions on both a state-action pair and the current hypotheses, (2) a theory generator that updates natural language hypotheses using Bayesian inference, and (3) a curiosity-driven reinforcement learning agent trained to collect evidence against the current hypotheses. Inspired by how humans, from children to scientists (Piaget, 1954; Gopnik and Meltzoff, 1997), actively update their internal world model by performing experiments, our agent’s exploration provides new evidence for hypothesis refinement, creating a virtuous cycle of improvement.
@@ 2.1 Problem statement
In such a setting, we study how to rapidly learn an LLM-based parametric model $\mathbb{P}^{LLM}$ of the transition function $\mathcal{T}$ (also called forward model) given a budget of experiences $T$ allowed to an agent. This parametric model outputs a probability distribution over all possible next states given a state-action pair. The natural approach from the RL literature to improve this forward model uses gradient-based updates (Ha and Schmidhuber, 2018; Hafner et al., 2020; Schrittwieser et al., 2020). However, the scale of modern LLMs makes such an approach particularly costly. LLMs’ natural in-context learning abilities open up an alternative approach: providing natural language hypotheses in the prompt to improve their predictions. This method is less costly, known as more sample efficient (Le Scao and Rush, 2021), and provides human-interpretable updates. Consequently, the parameter we seek to optimize is a set of hypotheses $\mathcal{H}\subseteq\mathcal{V}^{N}$ that, when given as input to the forward model (in addition to state-action pairs), maximizes the likelihood of the experienced next states.
@@ 2.2 WorldLLM
To improve the Statistician’s prediction as a forward model, the Scientist’s role is to search through the space of possible sets of hypotheses. To efficiently explore this space, we take inspiration from (Wang et al., 2024; Piriyakulkij et al., 2024) and use a Bayesian inference framework. In particular, we rely on another LLM $\mathbb{P}^{Sc}$ as the proposal distribution which generates a candidate set of hypotheses following the distribution $\mathbb{P}^{Sc}(\hat{H}_{t}^{i}|\mathcal{D}_{t},\hat{H}_{t})$ given a data set of collected transitions $\mathcal{D}_{t}$ and the current best hypotheses candidate $\hat{H_{t}}$. We assume a uniform prior (i.e., no particular prior) over the space of hypotheses. For this paper, we chose the Metropolis algorithm, a variant of the Metropolis-Hastings algorithm, which iteratively generates candidate hypotheses for $n_{s}teps$. The full algorithm is displayed in Algorithm 1 and more details are provided in Appendix B.
@@
Then, we introduce RL-based Experimenters. We propose to study three curiosity-based intrinsic rewards derived from $\mathbb{P}^{St}$ (Oudeyer and Kaplan, 2007). First, we introduce RL-LogP, where the Experimenter uses the prediction error (Schmidhuber, 1991a; Pathak et al., 2017; Burda et al., 2019a) as reward by receiving $r=-log\mathbb{P}^{St}(s^{\prime}|s,a,H_{t})$ for each collected transition $(s,a,s^{\prime})$. While usually efficient, this method is also known to suffer in stochastic environments where it fails to separate epistemic from aleatoric uncertainty (Burda et al., 2019b). Second, we propose RL-ALP, which leverages information gain over the prediction error (a form of Learning Progress) (Oudeyer and Kaplan, 2007; Lopes et al., 2012; Schmidhuber, 1991a) and rewards transitions on which the forward model improves. In particular, we use the absolute value of Learning Progress, which tracks both progress and forgetting in the forward model (Oudeyer and Kaplan, 2007; Baranes and Oudeyer, 2009):
@@
Finally, directly tracking Learning Progress at the transition level often leads to noisy estimations. We also propose RL-ALPEXP, a variant of RL-ALP which partitions the transition space into subspaces and computes ALP over each subspace, stabilizing the estimations. While existing works studied how to automatically infer such partitions (Baranes and Oudeyer, 2013), we here rely on hand-designed ones, which correspond to the partitions used by oracles. Further details are provided in Appendix A.
@@ 3.1 Experimental setup
To evaluate WorldLLM’s efficiency, we use the Playground-Text environment from (Gaven et al., 2024), a textual environment which produces textual observations and receives textual commands. This environment generates a room that contains multiple objects the agent can interact with by grasping or releasing them. The agent can combine two objects, if an interaction is possible, by releasing one onto the other, as illustrated in Figure 2 b. For instance, releasing water on a carrot seed transforms them into grown carrot. In our experiments, we only consider 4 types of objects: Water, Plants (seed or grown) that require water to grow, Small Herbivores (baby or grown) that require a grown plant to grow, and Big Herbivores (baby or grown) that require 2 grown plants to grow. We evaluate WorldLLM’s ability to discover, predict, and explain how to grow Plants, Small Herbivores, and Big Herbivores. We provide more details on the environment in Appendix A. For analysis, we gather the possible transitions into 6 different types: Standing when moving on an object, Holding 1 when grasping an object, Holding 2 when grasping two objects at the same time, Grow Plant when releasing water on a seed, Grow S. Herbivore when releasing a grown plant on a baby small herbivore and Grow B. Herbivore when releasing two grown plants on a baby big herbivore. These types are the ones used by RL-ALPEXP.
@@ 4.1 Building forward models
Building forward models is a long-established research direction (Ha and Schmidhuber, 2018; Moerland et al., 2023). In addition to producing more interpretable agents, forward models were particularly used in RL to improve sample efficiency, allowing the agent to train on imagined transitions or perform multi-step planning (Hafner et al., 2020; 2021; 2025; Schrittwieser et al., 2020). Parallel to this, learning a forward model has been used as an intrinsic reward to foster RL policies’ exploration (Schmidhuber, 1991b; Pathak et al., 2017; Burda et al., 2019a; Oudeyer and Kaplan, 2007). WorldLLM employs a similar curiosity-driven scheme where our RL-based Experimenters use reward signals from prior work to promote exploration and collect transitions that help improve the forward model. However, our method differs in the way the model is improved: instead of using a gradient-based approach to fine-tune the LLM-based forward model, as done in (Xiang et al., 2023), WorldLLM searches for natural language theories that improve the forward model when given in the prompt. While recent work studied how a forward model could leverage external textual information such as a manual explaining the environment (Zhang et al., 2024; Dainese et al., 2023), our approach iteratively generates and refines its own explanations.
@@ 4.2 Inductive reasoning for world modeling
The importance of theories in shaping one’s representations about the world was argued as central, in both children development and adults (Piaget, 1954; Gopnik, 1996) Building on this, the literature of theory-based RL has been advocating for augmenting artificial agents with inductive reasoning (Tsividis et al., 2021). In particular, (Tsividis et al., 2021; Tang et al., 2024; Li et al., 2024) proposed approaches based on LLMs that generate programs representing the theories about an environment’s dynamics. The use of programs allowed the authors to execute and compare how the proposed theories match empirical evidence. Parallel to this, (Piriyakulkij et al., 2024; Wang et al., 2024) proposed to generate theories using natural language. While the former constrained the generated theories to programs that can be easily evaluated in their environment, the latter used the generated theories to condition a program generator to simulate the environment’s dynamics. WorldLLM also builds on the theory-based RL literature and closely relates to (Piriyakulkij et al., 2024). Notably, WorldLLM also leverages Bayesian inference with an LLM as the proposal distribution to generate the theories. However, the major difference lies in how the produced theories are used: while (Piriyakulkij et al., 2024) constrain the format of theories such that they can be directly tested in the environment, our work focuses on giving these theories in another LLM’s prompt to improve its forward modeling capabilities. In other words, our work strongly differs in its perspective of producing theories that improve our LLM’s world modeling abilities instead of producing correct theories. Nonetheless, the Bayesian inference framework we used in our Scientist is largely inspired by (Piriyakulkij et al., 2024). Both our work and theirs also consider an active learning paradigm to choose which experiments to perform for theories’ refinement. While their approach computes exact Information Gain of all possible experiments, ours relies on curiosity-driven RL agents and Information Gain approximates as rewards to overcome the intractability of exact Information Gain in most environments.
@@ 5 Discussion
Despite promising results, our initial implementation of WorldLLM presents several limitations. First, the Playground-Text environment is structurally simple, and the natural ordering created by the most challenging transitions being longer and harder to predict is not expected in more complex environments—potentially reducing the effectiveness of RL-LogP. Evaluating WorldLLM in more complex environments is therefore a critical direction for future work (Ying et al., 2025).
@@
In addition to scaling the LLM, we argue that the emergence of such abstract hypotheses may result from other modifications to our implementation. A key advantage of Bayesian inference in this context is the ability to incorporate priors, which constrain the theory space. While a length-based prior might encourage abstract hypotheses, we believe richer priors—such as “error maps” that represent structured patterns of predictive failure, as described by Schulz (2012)—may be necessary to better replicate human-like theory refinement.
@@
Finally, our exploration strategy relied solely on simple curiosity-driven signals focused on poorly predicted transitions. Other intrinsic motivation mechanisms have been studied in both humans and artificial agents. A particularly important one is autotelic learning, where learners generate, select, and attempt to master self-defined goals (Steels, 2004; Colas et al., 2022). Studying the use of such more advanced curiosity signals could significantly improve the quality of collected transitions.
@@ Appendix A The Playground-Text Environment
We use the Playground-Text environment from (Gaven et al., 2024), itself being a textual adaptation of the Playground environment introduced by (Colas et al., 2020). This textual environment returns, at each time step, a description of its state and possible actions. In our experiments, we used a modified version of Playground-Text where the current state still describes the whole environment’s state, but the next state only describes how the performed action changed the environment (see Figure 6).